OpenStudy (anonymous):

please tell me where i went wrong:

OpenStudy (anonymous):

\[9^{-3}9^{\frac{ -3 }{ 2}}\] \[\frac{ 1 }{ 9^{3} } \times \frac{ 1 }{ 9^{\frac{ 3 }{ 2}}}\]

OpenStudy (anonymous):

okay so far?

OpenStudy (anonymous):

@nickhouraney

OpenStudy (anonymous):

oops got distracted with the other problem lol

OpenStudy (anonymous):

lol its ok, you can finish that first :)

OpenStudy (anonymous):

im having trouble seeing this is it 9^-3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i believe you are correct

OpenStudy (anonymous):

@lgbasallote ?

OpenStudy (lgbasallote):

you guys carry on. im just watching

OpenStudy (anonymous):

no watching lol

OpenStudy (anonymous):

okay but if we simplify further: \[\frac{ 1 }{ 9^{\frac{ 9 }{ 2 }} }\]? my book's answer is \[\frac{ 1 }{ 3^{9} }\]

OpenStudy (lgbasallote):

watching is fun. makes people get nervous >:))

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thats the answer!! D:

OpenStudy (anonymous):

please anyone?

OpenStudy (anonymous):

workin on it...

OpenStudy (anonymous):

If bases are same then can't we add the powers ??

OpenStudy (anonymous):

\[\large x^{-a} \times x^{-b} \implies x^{-a - b}\]

OpenStudy (anonymous):

well i got the answer but it appears i did it a little different

OpenStudy (anonymous):

thats what i did.. added the powers

OpenStudy (anonymous):

@nickhouraney how did you do it?

OpenStudy (lgbasallote):

you forgot to change to common denominators

OpenStudy (lgbasallote):

darn @waterineyes you were supposed to only watch and pressure @nickhouraney :p

OpenStudy (anonymous):

ya i was getting too nervous lol

OpenStudy (anonymous):

Sorry @lgbasallote

OpenStudy (anonymous):

Leave one comment here @ajprincess

OpenStudy (anonymous):

okay, so change denominators..

OpenStudy (anonymous):

how i did it was simplify it so i got 1/729 * 1/27

OpenStudy (anonymous):

@ajprincess write one comment here..

OpenStudy (anonymous):

but i won't have a calculator for such small numbers..

OpenStudy (anonymous):

@nickhouraney don't worry whole solution is coming right now..

OpenStudy (angela210793):